Medical command gives you the following orders for a rapid sequence intubation. You estimate the patient weighs at 175 pounds.

Amidate is 40mg/20ml

Your order for Amidate is 0.3 mg/kg.

What is the correct dose of Amidate for this patient ?
How many mL would you draw from this vial?
Answer #1

Ans) Amidate is 40mg/20ml

Order for Amidate is 0.3 mg/kg.

Weight: 175 pounds

175/2.205 = 79.3 kg

0.3 × 79.3 = 23.79 mg

40 mg = 20 ml

23.79 mg = x

x= 23.79 × 20/ 40 = 11.89 ~ 12 ml.

The correct dose of Amidate is 12ml.

12 ml will be drawn from this vial.